# A Comprehensive Evaluation of Large Language Models on Temporal Event Forecasting

### v1

@@ 4.3. Effects of RAG (RQ2)
4.3.1. Performance w.r.t. Various Retrieval Models
@@
4.3.2. Performance w.r.t. Various Retrieval Scopes
Table 6. Performance comparison w.r.t. applying different retrieval scopes. ”N/A” means ”Not Applicable”. $\ast$” indicates no historical input and it is falls into graph-only.
@@
Table 7. Performance comparison of using different retrieval models on the Text-only and Mixed settings.
Retriever Text-only Mixed
BM25 (Robertson et al., 2009) 0.2893 0.3129
Contriver (Izacard et al., 2021) 0.3015 0.3285
LlamaIndex (Liu, 2022) 0.3228 0.3539
4.3.3. Performance w.r.t. Varying Historical Length
The scope of history can be interpreted as the spatial dimensionality of the world, while the temporal dimensionality does also matter. Especially in the task of temporal event forecasting, properly setting the historical length, even with the utilization of retrieval models, may play a crucial role in effective forecasting. Driven by this hypothesis, we test our default model of ”retrieved history” in terms of different historical lengths, ranging from $\{3,7,15,30,90\}$ days. In other words, we constrain the retrieval models to search within the specified period of history. For example, three days of historical length means the retrieval models can only search from the events/articles in the past three days. We demonstrate the results in Figure 3. For graph-involved methods (Graph-only and Mixed), the performance first grows and then drops, with the historical length increasing. This phenomenon makes sense because when the given history is very short, the useful information for sound forecasting is likely to be in the farther past. Hence, no matter how powerful the retrieval model is, it will be bottle-necked by the limited accessible history. When the historical length increases, more information is brought to the view, resulting in increased performance. However, the expansion of history length comes along with more noisy information, which is why when the historical length exceeds some threshold, the performance starts dropping. Given its poor performance, such phenomenon does not appear in the Text-only setting. In this case, it may be because the text-based retrieval models constrain the effectiveness of the retrieval process, making the forecasting method insensitive to historical length.
@@
Figure 3. Performance comparison considering varying historical length for the model of ”retrieved history”.
4.4. Effects of Popularity Bias (RQ3)
@@
Several previous works, such as SeCoGD (Ma et al., 2023a), have identified that event forecasting suffers from severe popularity bias. To study this interesting problem, we group the testing atomic events into four clusters according to the sparsity (popularity) degree of the object entity  [^7] Alternative ways to measure the sparsity degree of an atomic event are using the occurrence frequency of subject entity or the average frequency of both subject and object. We leave these additional settings for future work. within each quadruple. As shown in Figure 4, each bar corresponds to one group of entities and its x-axis label denotes the occurrence frequency span. The height of the bar represents the number of atomic events whose object entities fall in the group, meanwhile, we separately annotate the number of atomic events in train/val/test sets for each group. The line denotes the number of entities (including both subject and object entities) in each group. Apparently, with a similar size of atomic events, the sparser groups have larger sets of entities while denser groups have fewer, exhibiting significant popularity bias (long-tail) phenomenon.
@@
Figure 4. The statistics of different sparsity (density) groups, where bars show the number of atomic events and lines indicate the number of entities in each group.
Figure 5. The breakdown performance comparison on different sparsity (density) groups, where the top sub-figure shows two denser (more popular) groups while the bottom sub-figure illustrates two sparser groups. The horizontal axis corresponds to multiple methods including both non-LLM method (LoGo) and LLM-based methods.
5. Related Work
@@
5.1. Temporal Event Forecasting
@@
Temporal event forecasting aims to predict future events based on the observed historical events. Different formulations have been raised for this task (Zhao, 2022). They can generally be grouped into three categories based on the event format: time series, unstructured textual events, and structured events.
@@
Some works model event occurrence or features using time series (Weidmann and Ward, 2010; Benjamin et al., 2023; Morstatter, 2021), but they fail to represent multi-relations among entities and multi-line natural of events. Several studies also explore the unstructured textual representation of temporal events, where each atomic event is generated from multi-documents in the form of summary (Gholipour Ghalandari et al., 2020) or phrases (Jiao et al., 2023) and is chained in temporal order. The natural language event description contains more fine-grained details but also leads to higher information complexity that impedes the downstream forecasting task formulation and performance. ForecastQA (Jin et al., 2021) formulates the event forecasting task in a question-answering manner with a retrieval database consisting of all documents that contain historical events. However, due to the lack of structure extracted events, the generation of QA pairs requires heavy human labeling and extensive domain knowledge.
@@
Efforts have also been exhausted in structured event forecasting. The major line of work represents temporal events using temporal knowledge graphs where each atomic event is a time-stamped link (Leetaru and Schrodt, 2013; O’brien, 2010). Extensive methods have been proposed for forecasting by aggregating the temporal and relational relation among entities (Jin et al., 2020; Li et al., 2021b; Park et al., 2022), or retrieving relevant historical events (Zhu et al., 2020; Sun et al., 2021; Li et al., 2021a), or modeling a continuous time development of events (Trivedi et al., 2017; Ding et al., 2021). Some methods have also tried to incorporate textual event information into TKG. Glean (Deng et al., 2020) and CMF (Deng et al., 2021) fuse textual embeddings into graph edges, SeCoGD (Ma et al., 2023a) uses textual topic modeling to disentangle subgraphs, and the dataset MidEast-TE and model LoGo (Ma et al., 2023b) leverages text clustering to construct complex event for forecasting with local and global contexts. However, all of them still conduct the forecast reasoning on graphs, while in our work, we investigate forecasting in a hybrid setting of leveraging both text and graph.
5.2. LLMs in Temporal Event Forecasting
@@
Generative Language Models (LM), especially the Large Language Models (LLMs) (OpenAI, 2023), have exhibited superior capability in language understanding and reasoning in various tasks and domains, such as science (Chen et al., 2023; Wang et al., 2023; Lu et al., 2024) and healthcare (Wu et al., 2023; Thirunavukarasu et al., 2023). Researchers have also conducted various studies on LLMs for temporal reasoning. One line of work focuses on temporal understanding where LLMs are tested for temporal event ordering or storyline understanding (Tan et al., 2023; Ning et al., 2020; Zhou et al., 2019; Zhang and Choi, 2021; Wang and Zhao, 2023). Compared to the understanding task, the forecasting task is generally of higher difficulty where the prediction targets do not appear in the input and therefore require the model to conduct necessary inferences. Several studies have leveraged LLMs for temporal event forecasting by converting the TKG formulation into text sequence and converting missing object prediction into next token prediction (Xu et al., 2023). GPT-NeoX-ICL (Lee et al., 2023) leverages in-context learning of LLMs and constructs prompts as a list of historical events each in quadruplet format. GENTKG (Liao et al., 2023) improves the selection of historical event inputs by a temporal logical rule-based retrieval strategy. LAMP (Shi et al., 2023) instead applies LLMs to perform abductive reasoning to assist the retrieval. However, all of these works only investigate LLMs on temporal event forecasting with structured graph event data. The evaluation of LLMs on textual temporal event forecasting remains explored, not to mention the hybrid setting of graph and text that we also study in this work.
6. Conclusion and Future Work
@@
Moving forward, to address the identified key challenges, there is a lot of work to be done, and here we would like to highlight two aspects. First, building larger, more accurate, and versatile benchmark datasets is pressing. Event forecasting is featured with highly domain-specific characteristics, therefore, it is hard to obtain a generalizable conclusion on small-scaled noisy datasets in one or a small number of domains, which is also the major limitation of this work. Second, focusing on task-level characteristics instead of arms racing on overall performance is necessary and more valuable. Event forecasting is a long-standing but slow-developing research area, one non-negligible reason of which is under exploration and poor understanding of problem settings.
Appendix A Appendix
A.1. prompts
In this section, we show the prompts designed for various modules. Table  9 and Table  10 show the prompts of news summarization and graph retrieval, respectively. As shown in Table 8, the red section represents the rule-based method, while the blue section represents the retrieval-based method.
Table 8. Prompt template of event forecasting
Table 9. Prompt template of summarization
You are an assistant to summarize news text with the following rules: 1. You need to generate a concise summary based on the news text. 2. Your response should only include the generated summary.
Table 10. Prompt template of graph retrieval

### v2

@@ 3.3 Effects of RAG (RQ2)
3.3.1 Performance w.r.t. Various Retrieval Models
@@
Table 9: Performance comparison of using different retrieval models on the Text-only and Mixed settings.
Retriever Text-only Mixed
BM25 [34] 0.2893 0.3129
Contriver [35] 0.3015 0.3285
LlamaIndex [36] 0.3228 0.3539
Table 10: Performance comparison w.r.t. applying different retrieval scopes. ”N/A” means ”Not Applicable”. $\ast$” indicates no historical input and it is falls into graph-only.
@@
Figure 4: Performance comparison considering varying historical length for the model of ”retrieved history”.
3.3.3 Performance w.r.t. Varying Historical Length
The scope of history can be interpreted as the spatial dimensionality of the world, while the temporal dimensionality does also matter. Especially in the task of temporal event forecasting, properly setting the historical length, even with the utilization of retrieval models, may play a crucial role in effective forecasting. Driven by this hypothesis, we test our default model of ”retrieved history” in terms of different historical lengths, ranging from $\{3,7,15,30,90\}$ days. In other words, we constrain the retrieval models to search within the specified period of history. For example, three days of historical length means the retrieval models can only search from the events/articles in the past three days. We demonstrate the results in Figure 4. For graph-involved methods (Graph-only and Mixed), the performance first grows and then drops, with the historical length increasing. This phenomenon makes sense because when the given history is very short, the useful information for sound forecasting is likely to be in the farther past. Hence, no matter how powerful the retrieval model is, it will be bottle-necked by the limited accessible history. When the historical length increases, more information is brought to the view, resulting in increased performance. However, the expansion of history length comes along with more noisy information, which is why when the historical length exceeds some threshold, the performance starts dropping. Given its poor performance, such a phenomenon does not appear in the Text-only setting. In this case, it may be because the text-based retrieval models constrain the effectiveness of the retrieval process, making the forecasting method insensitive to historical length.
Figure 5: The performance comparison on long-term forecasting, where the sub-figure shows the accuracy of the different ”retrieved history” models. The horizontal axis represents the time interval between the current timestamp and the query timestamp.
3.3.4 Performance w.r.t. Long-Term Forecasting
In temporal event forecasting, forecasting events over longer time horizons can be more challenging due to increased uncertainty, the influence of external factors, and the potential for unforeseen events. To evaluate the performance of LLM-based methods in long-term forecasting, we tested our default ”retrieved history” model to predict events occurring within a future time frame ranging from 1 to 14 days. For example, a time interval of 7 days signifies using historical events before timestamp t to predict events occurring at $t+7$. As shown in Figure 5, the performance of LLM-based methods, as expected, deteriorates as the temporal gap between future and current timestamps increases. Furthermore, the variation in prediction accuracy across different time points may result from the uneven distribution of events at various stages of development. In other words, recent events may not be highly related to the query event, meaning that short-term changes in the query event might not significantly influence the prediction. Additionally, the performance on the mixed setting is worse than the graph-only and text-only settings. This may be because the mixed retrieval method provides more fine-grained contextual information but also introduces a considerable amount of irrelevant noise.
Figure 6: The statistics of different sparsity (density) groups, where bars show the number of atomic events and lines indicate the number of entities in each group.
3.4 Effects of Popularity Bias (RQ3)
@@
Several previous works, such as SeCoGD [11], have identified that event forecasting suffers from severe popularity bias. To study this interesting problem, we group the testing atomic events into four clusters according to the sparsity (popularity) degree of the object entity  [^8] Alternative ways to measure the sparsity degree of an atomic event are using the occurrence frequency of subject entity or the average frequency of both subject and object. We leave these additional settings for future work. within each quadruple. As shown in Figure 6, each bar corresponds to one group of entities and its x-axis label denotes the occurrence frequency span. The height of the bar represents the number of atomic events whose object entities fall in the group, meanwhile, we separately annotate the number of atomic events in train/val/test sets for each group. The line denotes the number of entities (including both subject and object entities) in each group. Apparently, with a similar size of atomic events, the sparser groups have larger sets of entities while denser groups have fewer, exhibiting significant popularity bias (long-tail) phenomenon.
Figure 7: The breakdown performance comparison on different sparsity (density) groups, where the top sub-figure shows two denser (more popular) groups while the bottom sub-figure illustrates two sparser groups. The horizontal axis corresponds to multiple methods including both non-LLM method (LoGo) and LLM-based methods.
@@
Figure 8: Case Study: Examples of our methods that consider all three input formats, encompassing both rule-based and retrieved history.
3.5 Case Study
In this section, we conduct a qualitative analysis of our methods w.r.t. various input formats and histories. The details of these cases are presented in Figure 8 , from which we can observe several interesting findings. First, graph-based history is more effective at modeling temporal and relational patterns, but it also amplifies the impact of noise. For example, in the graph-only setting, rule-based history includes multiple historical events that provide rich historical information but may interfere with the reasoning capabilities of LLMs. Conversely, retrieved history is constructed by extracting triples, such as (Justice Ministry, Investigate crime corruption, Police, 2540). The connection between ”Police” and ”NSO Group tool Pegasus” enables LLMs to infer the correct answer. Second, news texts can provide richer contextual information for event forecasting, thereby aiding the reasoning process of LLMs. However, the inclusion of historical news texts in the input is limited by token constraints and the ability of LLMs to handle long texts. Compared to rule-based history, retrieval-augmented generation (RAG) can significantly reduce the token length of the input while retaining the most relevant historical information. Nonetheless, retrieved news texts based on semantic similarity do not necessarily reflect the development trends of events, which can lead to incorrect answers in the text-only setting when using retrieved history. This highlights the importance of more powerful retrieval mechanisms to improve the performance of text-based methods.
4 Related Work
@@
4.1 Temporal Event Forecasting
@@
Temporal event forecasting aims to predict future events based on the observed historical events. Different formulations have been raised for this task [1]. They can generally be grouped into three categories based on the event format: time series, unstructured textual events, and structured events.
@@
Some works model event occurrence or features using time series [37, 38, 39], but they fail to represent multi-relations among entities and multi-line natural of events. Several studies also explore the unstructured textual representation of temporal events, where each atomic event is generated from multi-documents in the form of summary [40] or phrases [41] and is chained in temporal order. Several studies have also investigated unstructured textual representations of temporal events, in which each atomic event is derived from multiple documents and expressed either as a summary[40], and then linked sequentially in temporal order. The natural language event description contains more fine-grained details but also leads to higher information complexity that impedes the downstream forecasting task formulation and performance. ForecastQA [17] formulates the event forecasting task in a question-answering manner with a retrieval database consisting of all documents that contain historical events. However, due to the lack of structure extracted events, the generation of QA pairs requires heavy human labeling and extensive domain knowledge.
@@
Efforts have also been exhausted in structured event forecasting. The major line of work represents temporal events using temporal knowledge graphs where each atomic event is a time-stamped link [2, 3]. Extensive methods have been proposed for event forecasting, including approaches that integrate temporal and relational dependencies among entities [12, 13, 14], retrieve relevant historical events [42, 43, 44], or model the continuous temporal evolution of events [45, 46]. RGCN [9] adopts a static knowledge graph to capture the relational patterns among entities. RENET [12] utilizes RGCN and the gated recurrent units (GRU) to model the evolution of the temporal knowledge graph over time for event forecasting. REGCN  [13] adopts ConvTransE [8] as the decoder to predict the event at the next time stamp. Some methods have also tried to incorporate textual event information into TKG. Considering the absence of specific and informative events,SeCoGD [11] employs textual topic modeling to disentangle subgraphs, leveraging the MidEast-TE dataset for event forecasting. In contrast, LoGo [6] harnesses text clustering to construct complex events for forecasting, incorporating both local and global contexts. In order to incorporate textual information into representation learning, Glean [4] and CMF [18] fuse textual embeddings in graph edges. However, all of them still conduct the forecast reasoning on graphs, while in our work, we investigate forecasting in a hybrid setting of leveraging both text and graph.
4.2 LLMs in Temporal Event Forecasting
@@
Generative Language Models (LM), especially the LLMs [47], have shown superior capability in language understanding and reasoning across a wide range of tasks and domains, including science[48, 49, 50] and healthcare [51, 52]. Researchers have also conducted various studies on LLMs for temporal reasoning. One line of work focuses on temporal understanding where LLMs are tested for temporal event ordering or storyline understanding [53, 54, 55, 56, 57]. Compared to the understanding task, the forecasting task is generally of higher difficulty where the prediction targets do not appear in the input and, therefore, require the model to conduct necessary inferences. Recent work has employed Large Language Models (LLMs) for temporal event forecasting by transforming Temporal Knowledge Graph (TKG) formulations into textual sequences, and recasting missing object prediction as a next-token prediction task [22]. For instance, GPT-NeoX-ICL[23] utilizes in-context learning by structuring prompts as a list of historical events, each represented in quadruplet format. Meanwhile, LAMP[58] applies LLM-based abductive reasoning to enhance retrieval in forecasting pipelines. GENTKG [24] learns temporal logical rules via temporal random walks and retrieves historical events by these rules to help LLMs reasoning on the TKG. However, all of these works only investigate LLMs on temporal event forecasting with structured graph event data. The evaluation of LLMs on textual temporal event forecasting remains explored, not to mention the hybrid setting of graph and text that we also study in this work.
5 Conclusion and Future Work
@@
Moving forward, to address the identified key challenges, there is a lot of work to be done, and here we would like to highlight two aspects. First, building larger, more accurate, and versatile benchmark datasets is pressing. Event forecasting is featured with highly domain-specific characteristics, therefore, it is hard to obtain a generalizable conclusion on small-scaled noisy datasets in one or a small number of domains, which is also the major limitation of this work. Second, integrating structural and textual information into the reasoning process of LLMs is crucial for improving the performance of temporal event forecasting. While fine-tuning and retrieval-augmented generation (RAG) show promise, they remain constrained by token length limitations and the inherent gap between structured graph data and unstructured textual data. Thus, exploring the alignment of graph representations with textual representations within the token embedding space of LLMs is a promising direction for future research. Finally, focusing on task-level characteristics instead of arms racing on overall performance is necessary and more valuable. Event forecasting is a long-standing but slow-developing research area, one non-negligible reason of which is under exploration and poor understanding of problem settings.
CRediT authorship contribution statement
He Chang: Conceptualization, Methodology, Validation, Writing-Original Draft. Chenchen Ye: Conceptualization, Investigation, Writing- Review & Editing. Zhulin Tao: Writing-Review & Editing, Project administration. Jie Wu: Investigation, DataCuration. Zhengmao Yang: Formalanalysis, DataCuration. Yunshan Ma: Conceptualization of this study, Resources, Writing- Original Draft, Supervision. Xianglin Huang: Resources, Writing-Review & Editing, Project administration. Tat-Seng Chua: Resources, Writing- Review & Editing, Project administration.